OpenStudy (anonymous):

Phillip charged $400 worth of goods on his credit card. On his first bill, he was not charged any interest, and he made a payment of $20. He then charged another $18 worth of goods. On his second bill a month later, he was charged 2% interest on his entire unpaid balance. How much interest was Phillip charged on his second bill?

OpenStudy (anonymous):

1st 400 2nd 400-20 =380 3rd 380+18=398 4th 2% of 398 = 7.96 5th 398+7.96 = amount he has to pay

OpenStudy (anonymous):

400 - 20 = 380 380 + 18 = 398 398 + ( 398 * .02 ) =405.96

OpenStudy (anonymous):

Thank you both very much for breaking it down for me :-)